The population in a town satisfies the logistic law.

dx/dt=x/100-x2/108

where t is the number of years after the 2019 population census. if the population x was 100,000 in 2019.determine

  1. population x as a function of time
  2. The year the population will be double
1
Expert's answer
2021-07-14T10:15:44-0400
"\dfrac{dx}{dt}=\dfrac{x}{100}-\dfrac{x^2}{10^8}"

"\int\dfrac{10^8dx}{10^6x-x^2}=\int dt"

"\int\dfrac{100dx}{x}+\int\dfrac{100dx}{10^6-x}=\int dt"

"100\ln|x|-100\ln|x-10^6|=t+100\ln C"

"\ln\dfrac{x}{|x-10^6|}=0.01t+\ln C"

"\dfrac{x}{|x-10^6|}=Ce^{0.01t}"

"x(0)=100000"


"\dfrac{10^5}{|10^5-10^6|}=Ce^{0.01(0)}"

"C=\dfrac{1}{9}"

"\dfrac{x}{|x-10^6|}=\dfrac{1}{9}e^{0.01t}"

1. If "x<1000000"


"9x=(10^6-x)e^{0.01t}"

"x(t)=\dfrac{10^6}{1+9e^{-0.01t}}"

2,

"x=2(100000)=2\times10^5"


"\dfrac{2\times10^5}{10^6-2\times10^5}=\dfrac{1}{9}e^{0.01t}"

"18=8e^{0.01t}"

"0.01t=\ln 2.25"

"t=100\ln2.25\ years"

"t=81.1\ years"